OpenStudy (anonymous):

Last problem... The square root of 2a+3 = a Answer is a=3, but I keep getting negative 3.

OpenStudy (anonymous):

\[\sqrt{2a+3}=a\]

OpenStudy (anonymous):

@Mertsj @satellite73

OpenStudy (mertsj):

Square both sides.

OpenStudy (anonymous):

2a+3=a^2

OpenStudy (mertsj):

And: \[a^2-2a-3=0\]

OpenStudy (mertsj):

Furthermore: \[(a-3)(a+1)=0\]

OpenStudy (anonymous):

AH, I see what you did. And -1 cannot be a solution, so 3 is the only correct answer. Thanks!

OpenStudy (mertsj):

yes. yw